OpenStudy (anonymous):

Prove: ( remember the correct format for a proof) 2^n+2^n=2^n+1

OpenStudy (anonymous):

\[2^n+2^n=2\times 2^n\] is a start

OpenStudy (anonymous):

(2)=1

OpenStudy (anonymous):

therefore, 2^n+1

OpenStudy (anonymous):

2^n+1+2+3+4?

OpenStudy (anonymous):

Is thar right?

OpenStudy (anonymous):

\[2^n+2^n=2\times 2^n=2^1\times 2^n=2^{n+1}\]